Write an equation of a line with a slope of 2/3 and y-intercept of 7. Write the equation in both slope intercept form and standard form.
Answer by mangopeeler07(462)

y=mx+b is the slope-intercept form of a line. m is the slope and b is the y intercept. So to write an equation of a line with a slope of 2/3 and y-intercept of 7, you would just plug in that slope, 2/3 for m and the y-intercept, 7 for b.

So y=mx+b becomes y=2/3x+7.